the sum of a number plus four times the number is 85. What is the number?
Answer by DrBeeee(684)

LET N = THE NUMBER
THEN WE HAVE
N + 4*N = 85 OR
5*N = 85 OR
N = 85/5 OR
N = 17
IS (17 + 4*17 = 85)?
IS (17 + 68 = 85)?
IS (85 = 85)? YES
ANSWER: THE NUMBER IS 17.